Determine the magnitude and direction of the force between two parallel wires 30 m long and 6.0 cm apart, each carrying 30A in the same direction. Determine the magnitude of the force between two parallel wires.
Answer #1
Concepts and reason

The concepts used in this problem are magnetic field due to a current carrying wire and the magnetic force. First, derive to expression for force between two wires using the concept of magnetic field due to a current carrying wire and the magnetic force. Then, calculate the force by substituting the values.

Fundamentals

Magnetic field due to a current carrying wire:

The expression for magnetic field due to a straight current carrying wire is:

B=μoI2πdB = \frac{{{\mu _{\rm{o}}}I}}{{2\pi d}}

Here, μo{\mu _{\rm{o}}} is the magnetic permeability, II is the current flowing in the wire and dd is the distance of point from the wire.

Magnetic force:

The magnetic force is perpendicular to the magnetic field and the direction of current. The magnitude of magnetic force is given by:

Fmag=L(I×B)Fmag=ILBsinθ\begin{array}{c}\\{{\vec F}_{{\rm{mag}}}} = L\left( {\vec I \times \vec B} \right)\\\\{F_{{\rm{mag}}}} = ILB\sin \theta \\\end{array}

Here, II is the current, vv is the velocity, BB is the magnetic field and θ\theta is the angle between current and the magnetic field.

The diagram for the given problem is:

Wire 1
Wire 2

The force on wire 2 is:

F=I2LBsinθF = {I_2}LB\sin \theta

Substitute 90o{90^{\rm{o}}} for θ\theta ,

F=I2LBsin90oF = {I_2}LB\sin {90^{\rm{o}}}

F=I2LBF = {I_2}LB …… (1)

The magnetic field at wire 2 due to current in wire 1 is:

B=μoI12πdB = \frac{{{\mu _{\rm{o}}}{I_1}}}{{2\pi d}} …… (2)

Substitute equation (2) in equation (1).

F=I2L(μoI12πd)F = {I_2}L\left( {\frac{{{\mu _{\rm{o}}}{I_1}}}{{2\pi d}}} \right)

F=μoLI1I22πdF = \frac{{{\mu _{\rm{o}}}L{I_1}{I_2}}}{{2\pi d}} …… (3)

The expression of force is:

F=μoLI1I22πdF = \frac{{{\mu _{\rm{o}}}L{I_1}{I_2}}}{{2\pi d}}

Substitute 30A30{\rm{ A}} for I1{I_1} , 30A30{\rm{ A}} for I2{I_2} , 30m30{\rm{ m}} for LL , 6cm6{\rm{ cm}} for dd and 4π×107H/m4\pi \times {10^{ - 7}}{\rm{ H/m}} for μo{\mu _{\rm{o}}} .

F=(4π×107H/m)(30m)(30A)(30A)2π(6cm)=(4π×107H/m)(30m)(30A)(30A)2π(6cm(102m1cm))=0.09N\begin{array}{c}\\F = \frac{{\left( {4\pi \times {{10}^{ - 7}}{\rm{ H/m}}} \right)\left( {30{\rm{ m}}} \right)\left( {30{\rm{ A}}} \right)\left( {30{\rm{ A}}} \right)}}{{2\pi \left( {6{\rm{ cm}}} \right)}}\\\\ = \frac{{\left( {4\pi \times {{10}^{ - 7}}{\rm{ H/m}}} \right)\left( {30{\rm{ m}}} \right)\left( {30{\rm{ A}}} \right)\left( {30{\rm{ A}}} \right)}}{{2\pi \left( {6{\rm{ cm}}\left( {\frac{{{{10}^{ - 2}}{\rm{ m}}}}{{1{\rm{ cm}}}}} \right)} \right)}}\\\\ = 0.09{\rm{ N}}\\\end{array}

Ans:

The force between two wires is 0.09N{\bf{0}}{\bf{.09 N}} .